Becker Brothers is the managing underwriter for a 1.80-million-share issue by Jay’s Hamburger Heaven. Becker Brothers is “handling” 11 percent of the issue. Its price is $26 per share, and the price to the public is $28.25. Becker also provides the market stabilization function. During the issuance, the market for the stock turns soft, and Becker is forced to purchase 45,000 shares in the open market at an average price of $27.00. It later sells the shares at an average value of $26.50. Compute Becker Brothers’ overall gain or loss from managing the issue. (Do not round intermediate calculations and round your answer to the nearest whole dollar.)
Answer #1

Becker brothers is handling 11 % of issue

No of shares handled by Becker brothers = 1.8 million x 11% = 198,000 shares

The profit or earnings on 198000 shares = 198000 * (28.25-26) = $445,500

The loss on sale of 50000 shares = 45000*(27-26.50) = $22500

overall profit from managing issue = $445500 - $22500 = 423,000

Overall gain = $423,000